List the members of these sets.

(a) {x | x is a real number such that x

2 = 5}

(b) {x | x is a positive integer less than 15}

(c) {x | x is the square of an integer and x < 200}

(d) {x | x is an integer such that x

2 = 5}


1
Expert's answer
2022-02-22T23:20:32-0500

Let us list the members of these sets.


(a) "A=\{x | x\text{ is a real number such that }x^2 = 5\}"

It follows that "A=\{-\sqrt{5},\sqrt{5}\}."


(b) "B=\{x | x\text{ is a positive integer less than 15}\}"

It follows that "B=\{1,2,3,4,5,6,7,8,9,10,11,12,13,14\}."


(c) "C=\{x | x\text{ is the square of an integer and }x < 200\}"

It follows that "C=\{0,1,4,9,16,25,36,49,64,81,100,121,144,169,196\}."


(d) "D=\{x | x\text{ is an integer such that }x^2 = 5\}"

Since the equation "x^2=5" has no integer roots, "D=\emptyset."